### v1

# Through the Lens of Human-Human Collaboration: A Configurable Research Platform for Exploring Human-Agent Collaboration

Abstract.
Intelligent systems have traditionally been designed as tools rather than collaborators, often lacking critical characteristics that collaboration partnerships require. Recent advances in large language model (LLM) agents open new opportunities for human-LLM-agent collaboration by enabling natural communication and various social and cognitive behaviors. Yet it remains unclear whether principles of computer-mediated collaboration established in HCI and CSCW persist, change, or fail when humans collaborate with LLM agents. To support systematic investigations of these questions, we introduce an open and configurable research platform for HCI researchers [^1] Our work’s homepage is available at https://human-agent-collab.hailab.io.. The platform’s modular design allows seamless adaptation of classic CSCW experiments and manipulation of theory-grounded interaction controls. We demonstrate the platform’s effectiveness and usability through two case studies: (1) re-implementing the classic human-human-collaboration task Shape Factory as a between-subject human-agent-collaboration experiment with 16 participants, and (2) a participatory cognitive walkthrough with five HCI researchers to refine workflows and interfaces for experiment setup and analysis.
human-agent collaboration, LLM agent, remote collaboration, research platform, open-science
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing HCI design and evaluation methods
Figure 1. The core architecture of our research platform, which encompasses four key modules: researcher interface allows researchers to configure experiment setup, monitor live sessions, and perform result analysis; participant interface enables participants to engage in the experiment with other participants and LLM agents; Agent Context Protocol ensures standardized agent integration with the experiment controller; experiment controller manages experiment states and actions updates, and synchronize all the interaction logs into the database.
1. Introduction
A persistent gap highlighted by the HCI and CSCW community is that intelligent systems are predominantly perceived as “tools,” rather than mutual collaborators with humans (Grudin, 2018; Shneiderman, 2022; Suchman, 1987). Despite their capacity to support humans in data analysis (Wang et al., 2019; Guo et al., 2024), design work (Oh et al., 2018; Lee et al., 2022), and decision-makings (Cai et al., 2019; Zhang et al., 2024), these systems typically lack the mutual awareness (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002), adaptivity (Allen et al., 1999; Amershi et al., 2019), accountability (Liao et al., 2020; Hoffman et al., 2023), and interdependence (Seeber et al., 2020; Zhang et al., 2021) that characterize effective partnerships. This deficiency often prevents users of intelligent systems to develop trust (Jian et al., 2000; Hoff and Bashir, 2015; Vereschak et al., 2021; Hancock et al., 2011), shared understanding (Van Rensburg et al., 2022; Clark and Brennan, 1991; Clark, 1996), or effective coordination (Schmidt and Bannon, 1992; Olson and Olson, 2000; Hinds and Bailey, 2003), which can further impede the quality of outcomes (Ashktorab et al., 2020; Zheng et al., 2023b), diminish human agency (Tapal et al., 2017; Kim et al., 2024), and undermine reliance in high-stakes situations (Parasuraman and Riley, 1997; Lee and See, 2004; Wang et al., 2020; Vaccaro et al., 2024).
The advent of LLMs has shifted the design space for intelligent systems (Brown et al., 2020; Bubeck et al., 2023). LLMs demonstrate the capacity to engage in fluid communication with humans (Sharma et al., 2023; Yang et al., 2024), while advanced prompting techniques further enable human-like reasoning and planning  (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2023a). Research on LLM agents (Park et al., 2023, 2024; Shanahan et al., 2023), which are LLM systems capable of exhibiting complex, human-like behaviors to solve tasks, shows these agents [^2] In this paper, we use ”LLM agent” and ”agent” interchangeably (Chen et al., 2025a). can yield distinct, believable cognitive and social behaviors, particularly through role-playing the provided persona profiles. These capabilities promise new paradigms of human-agent collaboration, where humans treat these LLM agents as mutual partners rather than mere tools. Yet the fundamental social and cognitive demands of effective collaboration remain (Seeber et al., 2020; Olson and Olson, 2013). The long-standing goal of treating machines as teammates requires moving beyond performance measurement to examine how social (Nass and Moon, 2000; Zhang et al., 2021), communicative (Liang et al., 2019; Zhang et al., 2023), and cognitive behaviors (Schelble et al., 2022; Ashktorab et al., 2020) co-evolve. To achieve flexible and mutual human-agent collaboration, prior work identified that we need interfaces (Amershi et al., 2019; Liao et al., 2020) and interaction paradigms (Wang et al., 2020; He et al., 2024) that help LLM agents communicate intent (Clark and Brennan, 1991; Liang et al., 2019), negotiate decisions (Straus and McGrath, 1994; Bos et al., 2002), manage shared context (Cramton, 2001; Gutwin and Greenberg, 2002; Kraut et al., 2003), and signal responsibility (Erickson and Kellogg, 2000; Hoffman et al., 2023; Liao et al., 2020) in ways humans can readily perceive and act upon.
To inform the design of interactions in the context of human-agent collaboration, our vision is that LLM agents should resemble remote human collaborators. Both LLM agents and remote human collaborators rely on structured channels such as text or voice, and both lack social cues (e.g., facial expressions, tone) that augment in-person interactions (Short et al., 1976; Walther, 1996; Olson and Olson, 2000). While this analogy is limited by differences in accountability, knowledge grounding, and identity persistence, it motivates designing interfaces and interactions for human–LLM-agent collaboration by drawing on decades of CSCW and HCI research on remote collaboration. Empirical and theoretical research has systematically identified key interaction variables in remote collaboration: communication modalities shape the richness of information exchange (Daft and Lengel, 1986; Straus, 1997; Bos et al., 2002), workspace awareness supports coordination (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002), and information transparency fosters trust and mutual understanding (Cramton, 2001; Hinds and Bailey, 2003). Recent studies confirmed that these foundational principles remain relevant as new agent capabilities emerge (Zhang et al., 2023; He et al., 2024; Lu et al., 2025). Grounding human-LLM agent interaction in this body of research allows HCI researchers to investigate which principles and challenges persist, which are shifted, and what new issues arise as boundaries between remote human collaborators and LLM agents blur.
A major barrier to pursuing this line of research is the lack of an open, configurable, and theory-driven research platform for systematically examining how diverse interaction designs influence human-agent collaboration. Conventional CSCW experimental paradigms, such as NeoCITIES (Schelble et al., 2022; McNeese et al., 2005), Shape Factory (Bos et al., 2004), DayTrader (Bos et al., 2002), and Passcode (Gero et al., 2020), feature distinct task protocols and interaction controls. Re-implementing them for human-agent collaboration requires considerable repetitive effort, and the resulting system still lacks reusability and adaptability across paradigms. This practice often creates obstacles to reproducibility and the cumulative extension of scientific findings. While general-purpose virtual lab platforms in social science provide infrastructure for group studies with humans, they lack agent integration protocols or CSCW-informed interface controls (Almaatouq et al., 2021; Chen et al., 2016). Existing agent-based frameworks emphasize multi-agent collaboration (Wu et al., 2024; Chen et al., 2023; Li et al., 2023), simulations (Park et al., 2023; Qian et al., 2023), or performance evaluation (Zhou et al., 2023; Leibo et al., 2021), rather than controlled experiments for investigating the interactions in human-agent collaboration. This lack of an extensible research platform impedes systematic, cumulative explorations and the goal of open science in the HCI community.
To address this gap, we present an open, configurable research platform for conducting reproducible, controlled experiments on human-LLM-agent collaboration. Our platform design draws inspiration from the Shape Factory experiment (Bos et al., 2004), a classic CSCW experiment for analyzing group dynamics between collocated and remote collaborators. The Shape Factory experiment supports rigorous manipulation of communication modality, awareness, social framing, and resource interdependence. Our platform enables HCI researchers to adapt classic experimental paradigms for human-agent collaboration research, such as strategic collaboration tasks (e.g., DayTrader (Bos et al., 2002)), collaborative decision-making tasks (e.g., Essay Ranking (Zheng et al., 2023b)), and collaborative task-solving (e.g., Passcode (Gero et al., 2020). The platform architecture features four key components, as shown in Fig 1:
  1. (1)
    researcher interface for manipulating parameters and interaction controls;
  2. (2)
    modularized and customizable participant interface that reflects those manipulations;
  3. (3)
    standardized agent context protocol ensuring consistent agent integration;
  4. (4)
    experiment controller serving as execution engine and logging system.
Critically, the platform allows systematic manipulation of theory-grounded interaction controls to investigate their impact on collaborative dynamics, such as communication modality (Straus, 1997; Bos et al., 2002), awareness dashboards (Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992), and social framing (Bos et al., 2004; Hinds and Bailey, 2003). Researchers can independently manipulate these controls and further customize the functionalities to meet more nuanced needs using the platform’s open-source code.
We demonstrate our platform’s utility through a two-part evaluation. First, we validate the usability and effectiveness of our platform for controlled experiments through two case studies that re-implemented the Shape Factory human-human collaboration experiment (six-person team) into a human-agent collaboration experiment (one person and five LLM agents as a team) (Bos et al., 2004). A total of 16 participants took part in a crossed, between-subjects design with counterbalancing in both case studies, where we manipulated a key interaction control in each. The goal was to confirm that our platform could faithfully re-implement a classic experimental paradigm and produce measurable differences in participant behaviors and outcomes. Analysis of behavioral logs and post-study surveys confirmed the platform’s ability to capture significant differences across conditions; for instance, varying communication modality led to distinct patterns in negotiation frequency and team performance, as well as substantial differences in perceived trust and workspace awareness, which is aligned with established findings in collaborative work. Second, we conducted a participatory cognitive walkthrough with five HCI researchers on human-AI collaboration to evaluate the researcher interface for experiment setup and data analysis. Their feedback informed iterative improvements to the workflow and interface design. Collectively, these evaluations validate our platform as a methodological foundation for the HCI community to build a systematic, evidence-based understanding of human-agent collaboration.
2. Related Work
We group and discuss relevant literature into four bodies of work: (1) the shift in HCI from AI as a tool to AI as a partner, (2) how LLMs change the collaboration landscape and what still remains unsolved, (3) CSCW research on remote collaboration that offers a lens for human-LLM agent collaboration, and (4) experimental paradigms and platforms for collaboration studies, and why a theory-driven, extensible platform for human–agent experiments is still missing but of significant importance to the HCI and CSCW community.
2.1. Human-AI Collaboration: From Tools to Partners
Intelligent systems have often been designed and perceived as tools, rather than partners that share goals, coordinate actions, and carry responsibility with people (Grudin, 2018; Shneiderman, 2022; Suchman, 1987). Research on human-AI collaboration highlights several gaps that impede intelligent systems from establishing partnerships with humans.
Effective partnerships depend on awareness (i.e., the ability to know who is doing what, where, and when (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002)) and on common ground (i.e., a shared understanding of context and references (Clark and Brennan, 1991) in collaborative systems). When systems obscure state, history, and intent, collaborators have to guess and repair misunderstandings, raising coordination cost and slows progress (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002; Clark and Brennan, 1991; Schmidt and Bannon, 1992; Clark, 1996). Besides, adaptivity and initiative are critical pillars to support mixed-initiative in collaboration, where a system should be able to decide when to act, when to wait, and how to make its moves understandable (Allen et al., 1999; Amershi et al., 2019).
Many intelligent systems, particularly before the emergence of LLMs, only offered template-based dialog or single-shot commands and failed to adapt to users’ evolving goals nor proactively offer the right help at the right time (Amershi et al., 2019). In addition, accountability was also underscored to allow human collaborators to ask a system why a recommendation was made, solicit supporting evidence, and examine uncertainty; otherwise, users may experience mis-calibrated trust and reliance (Liao et al., 2020; Hoffman et al., 2023; Jian et al., 2000; Lee and See, 2004; Hancock et al., 2011). Subsequently, trust mis-calibration often will result in poorer joint outcomes and reluctance to use systems in higher-stakes work (Parasuraman and Riley, 1997; Wang et al., 2020; Vaccaro et al., 2024). Moreover, collaborative outcomes highly depend on interdependence, where team members should engage with each other to negotiate roles and divide work (Seeber et al., 2020). Many work, however, treat systems as isolated predictors or recommenders and leave little room for negotiating and clarifying labor distribution, which will weaken shared understanding (Van Rensburg et al., 2022; Clark, 1996) and make coordination inaccessible (Schmidt and Bannon, 1992). When intelligent systems cannot fulfill the need as a partner and are instead recognized as tools, studies reported lower joint quality (Ashktorab et al., 2020; Zheng et al., 2023b) and reduced sense of agency and engagement (Tapal et al., 2017; Kim et al., 2024).
2.2. The New Landscape of Human-AI Collaboration in the Era of Large Language Model-based Agent
The emergence of LLMs sheds light on new opportunities for the landscape of LLM-based intelligent systems as mutual collaborators of humans. In particular, LLMs enable natural language as a shared medium that allows users to state goals, constraints, and references, while LLM systems can ask clarifying questions and propose plans both in free-form language (Brown et al., 2020; Bubeck et al., 2023; Sharma et al., 2023; Yang et al., 2024). This capability shifts interaction beyond rigid command structures into fluid, open-ended conversations. Furthermore, advanced prompting techniques, such as chain-of-thought (Wei et al., 2022; Kojima et al., 2022), enable LLMs to perform step-wise reasoning and planning for their predictions. These techniques allow visible forms of system accountability and initiative that people can read and negotiate.
Building on these technical achievements, researchers have developed “LLM role-playing agents,” which adopt personas and goals to simulate complex social and cognitive behaviors over extended interactions (Park et al., 2023, 2024; Shanahan et al., 2023). Applications are rapidly expanding: LLM agents have been explored as human surrogates for automated UX testings (Lu et al., 2025), and as tools for simulating populations and replicating classic findings in social science (Argyle et al., 2023; Aher et al., 2023).
Despite these advances of LLM agents, the fundamental social and cognitive demands for effective human-agent collaboration remain. The key lies in the design of interfaces (Amershi et al., 2019; Liao et al., 2020) and interaction paradigms (Wang et al., 2020; He et al., 2024) between humans and LLM agents to establish the collaboration needs identified in classic CSCW research. However, HCI researchers currently lack a dedicated research platform that allow systematic mapping of classic collaboration insights onto the interface features of LLM agents. Such a platform is essential for investigating how collaboration dynamics unfold in this new era of human–AI partnerships.
2.3. Human-Human Remote Collaboration as an Analytical Lens of Exploring Human-Agent Collaboration
LLM agents can be viewed as partially analogous to remote human collaborators: both rely on structured and explicit information channels, such as text or voice, and lack access to non-verbal cues (e.g., facial expressions, tone) that shape face-to-face interaction (Short et al., 1976; Walther, 1996; Olson and Olson, 2000). Early CSCW research on remote collaboration has identified key challenges and interaction variables that shape distributed teamwork. With respect to information channels, for instance, Media Richness Theory explains why low-bandwidth channels carry less contextual information and make negotiation harder (Daft and Lengel, 1986), with several experiments demonstrating that “leaner” media like text are less effective than “richer” media like video for conveying ambiguity and building consensus (Short et al., 1976; Straus, 1997). These results are critical to humans-agent collaboration because the interactions primarily happened via text. Interface cues such as typing indicators, meta messages, could enrich these channels and influence how users interpret agents’ intent and actions (Erickson and Kellogg, 2000; Iftikhar et al., 2023).
Another recurring theme is the “mutual knowledge problem”, which arises when collaborators lack “common ground” (Clark and Brennan, 1991). Without shared references and assumptions, people struggle to know what their partners know, leading to misunderstandings and flawed assumptions (Cramton, 2001). Interfaces that add timely acknowledgments and visible states of collaborators could reduce these failures (Kraut et al., 2003). Research on workspace awareness further strengthened this idea by defining cues about presence, activity, and intent that help people predict collaborators’ actions (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002). For LLM agents, this can be achieved by asking agents to generate future plans with rationales that could be made visible to human collaborators through interface designs. Task structure and coupling also matter. Olson and Olson (2000) showed why “distance matters”, highlighting how timing, handoffs, and availability complicate remote collaboration. Hinds and Bailey (2003) also analyzes how distributed teams coordinate under variable latency and partial information.
In particular, a variety of lab-based experiments were designed to systematically analyze the impact of different interaction variables in a controlled setting. For instance, in DayTrader, researchers manipulate the communication medium during negotiation in a social-dilemma setting and identified that the choice of medium shifts trust formation and deal quality (Bos et al., 2002). Bos et al. (2004) further illustrated trading biases and subgroup formation caused by the differences between in-group (collocated partners who can talk face-to-face) and out-group (remote collaborators) in the Shape Factory experiment. NeoCITIES imposes time pressure and role interdependence in a shared information space (McNeese et al., 2005; Schelble et al., 2022), whereas Passcode examines how teams form and update shared mental models during problem solving when only partial clues are available (Gero et al., 2020). Despite LLM agents may differ from human remote collaborators in how they explain their decisions and how stable their identity recognition and memory are, research in the explainability of AI has pointed to feasible approaches by enabling provenance, disclosure of uncertainties, and human-understandable rationales given LLMs’ strong reasoning and text generation capabilities (Liao et al., 2020; Hoffman et al., 2023).
Our research platform draws on this profound line of CSCW work by enabling the manipulation of different interaction variables on the interface designs through the re-implementation of conventional experimental paradigms, but in the context of human-agent collaboration.
2.4. Research-Focused Toolkits and Platforms for Collaboration Studies
Figure 2. Existing research-focused platforms for collaboration studies. From left to right, the top row: Share (Greenberg, 1990), GroupKit (Roseman and Greenberg, 1996), LabintheWild (Reinecke and Gajos, 2015), Empirica (Almaatouq et al., 2021); the bottom row: oTree (Chen et al., 2016), nodeGame (Balietti, 2017), Overcooked (Carroll et al., 2019), Pairit (Ju and Aral, 2025).
A growing number of research toolkits or platforms were designed to facilitate collaboration studies (Fig 2). We group them into four families: groupware/toolkits, general-purpose platforms, game-based research environments, along with newer human–AI collaboration systems and agent frameworks.
Early groupware toolkits in the HCI community provided fundamental building blocks for synchronous, distributed collaborative tasks between humans, which focused on enabling shared views, real-time session management, and widgets for customization, as seen in Share (Greenberg, 1990) and GroupKit (Roseman and Greenberg, 1996; Greenberg and Roseman, 1999). Researchers reflected on the evaluation of such toolkits with strategies outlined for evidence, generality, and adoption (Ledo et al., 2018). There exist other general-purpose online lab platforms that aim to lower the barrier for synchronous studies at scale. For instance, LabintheWild (Reinecke and Gajos, 2015) supports public-facing HCI experiments with built-in participant recruitment support and data analytics. Empirica (Almaatouq et al., 2021) provides real-time grouping, messaging, and logging for interactive tasks; oTree (Chen et al., 2016) and nodeGame (Balietti, 2017) implement economic games and behavioral experiments with session management and data capture. Nevertheless, these platforms only focus on human studies with no support for agent integration, and do not support manipulation of collaboration variables through interface customizations. In the HCI community, a number of studies also explored the usability of game environments as research platforms. Such game-based environments emphasize tightly-coupled coordination and performance, such as the real-time collaborative cooking game Overcooked (Carroll et al., 2019); researchers also leveraged Left 4 Dead 2 to compare perceptions of human versus NPC teammates during cooperative play (Wehbe et al., 2017).
On the other hand, a large number of research platforms were proposed for developing and evaluating LLM agents. Multi-agent frameworks compose LLM agents and tools for task automation or agent–agent interaction, such as AutoGen (Wu et al., 2024), AgentVerse (Chen et al., 2023), and CAMEL (Li et al., 2023). Another family of agent-based frameworks, such as WebArena (Zhou et al., 2023) and Melting Pot (Leibo et al., 2021), was designed to systematically assess agent performances and social behaviors in a simulated environment. Moreover, agent-based social simulacra demonstrate LLM agents’ capacity to model social and cognitive behaviors at scale in a sandbox environment (Park et al., 2023, 2024) or support large-scale simulation with thousands of LLM agents interacting in a spatial or societal environment (Piao et al., 2025). Nevertheless, these agent-oriented frameworks and simulation platforms emphasize agent performance, simulation fidelity, or agent–agent dynamics, not for human-agent interactions.
Our work fills the critical gap between the rich theoretical heritage of CSCW and the practical needs of modern human-agent interaction research. Existing platforms either emphasize human collaboration studies with no support for agent integration or, on the other hand, focus on multi-agent interactions and simulations with no human engagement; besides, there is a lack of research platforms that provide researchers the rigorous controls needed for human-agent collaboration research. One recently proposed field-experiment platform called Pairit shares a similar conceptual idea with our platform, in which researchers compare humans’ performance in a human-human setting and a human-agent setting for a co-edit task (Ju and Aral, 2025), but this platform lacks customizability for different experimental paradigms and interface designs. Unlike existing platforms, our proposed platform is explicitly designed for the manipulation of theory-driven collaboration variables, and it is extensible and configurable for HCI researchers to re-implement, adapt, and invent new paradigms beyond conventional experimental paradigms used in human collaboration research. More importantly, our platform is purpose-built for human-agent collaboration, with a standardized agent integration protocol for agent integration and a data logging module for collecting both human behavior and agent reasoning.
3. System Design
To address the need for reproducible and extensible research on human-agent collaboration, we designed and implemented an open-source, fully configurable research platform. It serves as a flexible infrastructure for controlled experiments by balancing ease-of-use with customizability. The platform’s architecture encompasses four components: Participant Interface for user interaction, Researcher Interface for experiment management, Experiment Controller that serves as the backend engine, and an Agent Context Protocol (ACP) to standardize agent integration.
Researchers can adapt our platform for different experimental paradigms in two ways: through a declarative Experiment Configuration Language (ECL) (example provided in Appendix B) that defines all entities, actions, and interface modules, or by directly extending the open-source codebase. This dual approach supports seamless adaptation of existing paradigms, such as social dilemma games (Bos et al., 2002) and collaborative decision-making tasks (Zheng et al., 2023b), and the development of entirely new ones. Researchers can adjust the experiment parameters and manipulate the interaction controls for different conditions through the researcher interface. Our design is guided by four principles derived from foundational CSCW and HCI research: generality and extensibility, modularity, theory-grounded customizability, and standardized agent integration.
We address open-science needs by releasing source code (upon acceptance) and by supporting the creation, validation, and sharing of configuration templates, expecting our platform to support the broader research community in human-agent collaboration research.
3.1. Design Principles and Rationale
The platform targets synchronous or near-synchronous, interface-mediated collaboration among multiple participants that may be all human, all LLM agents, or mixed human-agent teams. Four principles guide the design of our research platform, each of which is grounded in foundational HCI work.
The first principle is generality and extensibility across experimental paradigms. Instead of hard-coding the platform’s functionalities for a single experiment, we abstract task, resource, and participant models to a configuration framework that separates classes from instances. The configuration language was thus designed to adhere to this framework. With the configuration language, researchers can either re-use and customize existing classes or instantiate new classes for experiment-specific artifacts and visual elements, and define their custom attributes correspondingly. This object-centered design supports resource negotiation and exchange, information sharing, and collaborative decision making within a unified architecture. As specified in Bos et al. (2004), most experiment paradigms are “not intended to be a full simulation of real-world collaboration,” but focus on reproducing certain collaborative dynamics.
The second principle is the modularization of interfaces and logic. Modularization has been repeatedly identified by prior research as essential for adaptable and replicable experimental platforms (Greenberg and Roseman, 1999; Gutwin and Greenberg, 2002). In our platform, the participant and researcher interfaces are composed of independent modules whose visibility and parameters. The participant interface consists of separate modules for private states, social interactions, task operations, and shared awareness that can be independently toggled and customized. The researcher interface consists of four primary modules: experiment selection, session configuration, real-time monitoring, and post-study data analysis.
The third principle is the theory-grounded customizability of interaction variables. Currently, the platform organizes supported interaction variables into four layers drawn from CSCW and HCI research: information flow, action structure, social framing, and system responsiveness. Each variable is bound to dedicated interface design options or controller specifications where researchers can independently manipulate it; for instance, they can manipulate the communication channels to enable private messages, broadcast messages, or disable messaging for all participants. The specific interface modules and controller settings that correspond to each layer are detailed in Section 3.2.
The fourth principle is native, standardized support for agent integration to ensure experimental validity and reproducibility. We designed a rigorous ACP to standardize agent integration for consistency and reproducibility. The protocol serves two high-level purposes: aligning agents’ perceptions and interactions with the environment to mimic the same experimental constraints as remote human participants, and maintaining consistency of these constraints across different agents, regardless of agent architectures. The protocol standardizes agents’ access to the environment, action spaces, communication channels, and the structure of all data sent to and from the agents, which also supports comparative evaluation of different agent architectures natively.
3.2. Architecture Overview
The platform architecture follows a client–server model (Figure 1). Participant interfaces, ACP, and the researcher interface connect to an experiment controller over a real-time channel. In particular, ACP ensures that agents can periodically perceive the experiment states and interact with the experiment environment in a consistent way. The controller manages experiment state, validates and commits actions, and synchronizes all logs with the data storage. Here, we present design and implementation details of each component.
3.2.1. Participant Interface
Figure 3. An example of the participant interface for the Shape Factory experiment (Bos et al., 2004). The interface consists of five fully configurable modules: My Status displays participant’s private status, My Actions displays permitted actions, My Tasks displays the tasks to be completed, Social provides available interaction channels, and Information Dashboard shows other participants’ real-time information. The timer of the experiment is shown at the top.
The participant interface uses a three-column layout that generalizes across common experimental paradigms by separating visual modules for private state, social interaction, task operations, and shared awareness. An example participant interface for the Shape Factory experiment, which is used in the case study in Section 4.1, is shown in Figure 3.
The left column presents private state and individual actions. The My Status module shows individual resources and role-dependent capacities that are not necessarily visible to others, which could support studies of interdependence and strategic planning. The My Action module provides affordances for private actions, such as producing a resource or making an independent decision. The My Task module displays individual or collaborative goals and fulfillment status, which could support shared plans and reduce coordination overhead by making the task structure explicit.
The middle column contains the Social module for all participant interactions. This is supported by fruitful CSCW research that has underscored the importance of communication and negotiation affordances for studying trust, information grounding, and collaboration effectiveness (Straus, 1997; Bos et al., 2002; Hinds and Bailey, 2003). Here, a Trading page supports resource exchange and can be adaptive to broader collaborative decision-making activities. A separate Chat page supports text-based communication between participants. Researchers can configure trading with free-form or protocol-bounded offers to study negotiation dynamics and coalition formation (Poole et al., 1991). For the chat functionalities, we implemented three variations: private messages, group messages, or be disabled entirely. LLM agents can interact with human participants through trading and messaging by following the instructions and action spaces specified in ACP. The right column is a fully configurable shared Information Dashboard. The dashboard makes workspace awareness (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002) a customizable visual module; researchers can select what information to disclose, to whom, and at what level of granularity.
While this three-column layout serves as a default, the underlying modules are independent components that can be rearranged, enabled, or disabled through the Experiment Configuration Language for different experimental paradigms.
3.2.2. Researcher Interface
The researcher interface provides direct controls for researchers to design, execute, monitor, and analyze experiments for human-agent collaboration research. The final design, as shown in Fig 4, was refined through a participatory cognitive walkthrough (Polson et al., 1992) with five HCI researchers in human-AI collaboration to solicit design and interaction feedback and align the workflow with their needs (details in Section 4.2). The researcher interface organizes functionalities into three pages: experiment setup, real-time monitoring, and post-study analysis. For experiment setup, the interface guides researchers through four primary steps: experiment selection, parameter configuration, interaction control, and participant registration, as shown in Figure 4. In addition, researchers can customize participants’ information, including the persona profile for LLM agents, without editing the code.
The Experiment Selection module allows a researcher to load a saved session or create a new one. Once the session is created, it provides a list of experimental paradigms, each with a text description and a simplified illustration, so that researchers can become familiar with the purpose of different experimental paradigms, even if they do not know any of the experiments beforehand. Researchers can also upload a custom paradigm defined in an ECL file.
Then, the Session Configuration module offers researchers comprehensive control over (1) experiment parameters, such as initial resource allocation and experiment constraints; (2) interaction controls, such as communication channels and information dashboard (details in Section 3.3); (3) agent design, including parameters like response latency; (4) participant registration for both humans and agents.
Once the session is live, the Real-time Monitoring module allows the researcher to observe participant interactions and progress. After the session is completed, the Result Analysis module provides summary statistics and visualizations of participant performance and behavior. Researchers can view aggregate results or filter by participant, and can export raw event logs for further analysis.
Figure 4. The workflow of HCI researchers for study setup through the researcher interface, which includes five steps: experiment selection, parameter configuration, interaction control, participant registration, and participant customization (optional).
Researcher Workflow
An overview of HCI researchers’ workflow is illustrated in Figure 4. For the study setup, researchers begin by selecting whether they want to load an existing session or a new session; if the latter, they first need to choose the experimental paradigms that best suit their research interests. The system also enables researchers to upload a customized experiment configuration template, using the language specified in Section 3.2.3, if the provided ones do not meet their needs. Then, researchers will proceed to the session configuration module to customize interaction controls and agent-specific designs. The horizontal scrolling function of the researcher interface allows researchers to iteratively adjust the parameters in accordance with the experimental procedures or manipulate the interaction controls with different sets of parameters, which is highlighted by participants in the cognitive walkthrough study. After completing all the configurations, researchers conclude the setup with participant registration to decide the number of human and agent participants involved.
3.2.3. Experiment Configuration Language (ECL)
The Experiment Configuration Language (ECL) is a declarative language designed to give researchers fine-grained control over experimental paradigms without modifying the platform’s source code.
The ECL is specified in a structured format by defining its components through a set of core primitives: objects, actions, policies, and views. The language is organized around these four fundamental concepts. Objects represent the primary entities within the experiment, such as resources (e.g., money, parts), roles (e.g., buyer, seller), or abstract concepts like team scores. Each object is defined by a set of attributes, such as initial_valueinitial\_value or visibilityvisibility. Actions define the operations available to participants, specifying the actor, the costs an action incurs, and its effects on the state of the system’s objects. Policies are logical constraints that are specified by different experimental paradigms. These range from preconditions for specific actions (e.g., requiring sufficient resources) to global rules like time limits or winning conditions. Finally, Views separates the system’s internal state from the presentation layer by defining which objects and attributes are rendered in specific modules of the participant interface. This separation is critical, as it allows researchers to manipulate information visibility without altering the experiment’s core logic.
A concrete example of ECL for the Shape Factory is provided in Appendix B. In this example, $Money$ and $Shape$ are defined as objectsobjects. the $ProduceAction$ links the costs and effects for producing a shape, and is governed by a policy that validates the actor’s resources. The $Views$ declares what objects and states are visible in each visual module. During setup, the Experiment Controller parses an ECL file to dynamically instantiate the environment, enforce the specified policies, and render the correct views for each participant. The experiment controller also validates the configurations and reports conflicts when a new configuration template is uploaded. This approach allows researchers to create, share, and modify experimental paradigm designs in a transparent and reproducible manner.
3.2.4. Agent Context Protocol
Figure 5. The agent context protocol (ACP) provides a standardized channel between the experiment controller and LLM agents, regardless of agent architecture designs. ACP includes four modules: Experiment Rules (task setup and constraints), Experiment Specifications (participant roles and resources), Permitted Actions with Schemas (structured formats for communication and trades), and Private and Public States (information visible to participants and agents).
Inspired by the emerging standard of model context protocol [^3] https://modelcontextprotocol.io/ that connects LLMs with external tools and resources, we designed the ACP to standardize communication between LLM agents and the experiment environment (Figure 5). ACP aims to establish a standardized channel for perception, information transfer, and interaction between LLM agents and the experiment environment to ensure robust agent integration regardless of agent architecture designs. As a result, ACP enforces that LLM agents’ behaviors are aligned with the experimental rules and valid action spaces. More importantly, agents perceive the experiment and act within it under the same constraints and through the same structured channels as a remote human participant. This parity is essential for the validity of controlled, comparative studies.
At initialization, the experiment controller (Section 3.2.5) constructs an experiment-specific context for agents from the experiment configurations and the interaction conditions specified by researchers. The context includes participant IDs, groups, private and public states, permitted actions with schemas, communication channels, perception interval, and experiment rules. The perception interval is a time parameter for determining how often the experiment controller issues periodic state summaries of the current situation to the agents. The experiment controller will validate that the information in the state summaries is aligned with the private and public states specified in the agent context.
During a session, the state summary generated by the experiment controller also carries the current timestamp and remaining time in the session. Agents are expected to take the state summary as input and generate a response of actions, sometimes with rationales if they were asked to do so. The controller validates responses against the schemas and policies, returns an error message to the agent if an action was identified as invalid, and asks the agent to re-generate the response. Once the actions are validated, the controller commits these actions atomically, updates the environment state, and logs the whole interaction.
3.2.5. Experiment Controller
The experiment controller is the execution engine of the platform. During experiment setup, the controller creates a session instance, compiles experiment configurations, binds variables to the conditions specified by the researcher, and renders the user interfaces. During a live session, the controller manages the timer, validates actions and interactions for participants, updates private and shared state, and tracks all behavioral logs for post-hoc analysis. If an LLM agent is registered, the controller instantiates an agent context, as described above. At the same time, human participants can log in to the participant interface using their session ID and participant ID. The controller exhaustively tracks all the actions, including trades, messages, resource production and allocations, order submissions, as well as agent inputs and outputs.
3.3. Configurable Interaction Control for Collaboration
We implement four layers of configurable interaction variables that connect findings in classic collaboration research to interaction controls of the research platform. Among these controls, some lead to interface design variations and information visibility to participants, while others may impact the experiment constraints and agent behaviors. Through the researcher interface, users can configure controlled comparative experiments to analyze how different interaction controls impact collaboration dynamics. Users can also define and customize the controls through ECL.
Information Flow
This layer focuses on the type and visibility of information that is accessible to human participants and LLM agents. Prior research establishes that media characteristics shape mutual understanding (Clark and Brennan, 1991; Straus, 1997), awareness of others’ actions supports coordination (Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992), and information control affects strategic behavior in distributed teams (Cramton, 2001; Olson and Olson, 2000). We implement two primary controls with respect to these conceptual findings. The Information Dashboard can be enabled or disabled to show real-time, shared information to all the participants, such as each other’s trades, recent orders, inventories, and order fulfillment status. Researchers can further customize information visibility based on participant groups (e.g., all, human-only, agent-only, etc.), change the information update frequency, and adjust the granularity of information (i.e., real values versus text summary). The Communication Transparency can be configured for strictly private (one-to-one), group-chat, or disabled entirely. This lets researchers examine how information sharing and message visibility influence negotiation dynamics and group behavior, and may further customize this control with additional turn-taking and message length limits.
Action Structure
This layer defines the types of actions available, how they can be conducted, and the operational constraints that shape interdependence. Prior work links action coupling, protocols, and constraints to coordination cost, convergence, and coalition dynamics (Olson and Olson, 2000; Bos et al., 2004). The platform provides control over the Negotiation Protocol, which allows the trading interface to be configured for either strict accept-or-reject protocols or for open-ended negotiation with counter-offers. Researchers can also set constraints such as price limits or trade frequency. Researchers can also manage Concurrent Actions by either allowing participants to make multiple simultaneous trade offers, or participants must resolve one offer before making another
Social Framing
This layer captures how identity, roles, and goals are presented to participants. Such framing is known to shape trust, subgroup formation, and role negotiation in distributed teams (Bos et al., 2004; Hinds and Bailey, 2003; Straus and McGrath, 1994). Researchers can manipulate Agent Identity by customizing the display names of LLM agents (e.g., ”teammate,” ”assistant”) and controlling the visibility of their personas or group affiliations. The platform also supports configurable Social Cues, such as color schemes or group identifiers, to explore the effects of in-group and out-group framing on collaboration.
Agent Responsiveness
This layer controls the timing and feedback mechanisms of LLM agents. Timing of responses can affect perceived competence and availability in remote collaboration (Olson and Olson, 2000), while adaptive communication and clear feedback help users recover from errors and calibrate reliance (Amershi et al., 2019). Key controls include Agent Response Latency, which can be set to be immediate or include a fixed or variable delay. Researchers can also customize a ”typing” or ”thinking” indicator. Agent Adaptive Feedback configures whether LLM agents use static response styles or dynamically adapt their strategy based on interaction history, specified through their persona profile. Finally, researchers have control over agents’ Explanations and Error Handling by deciding whether agents provide rationales proactively or on-demand. This control is extensible to support advanced agent reasoning architectures, such as the fast-and-slow thinking systems used in recent agent research (Lu et al., 2025).
3.4. Adaptivity Across Classic Experiments
The combination of the Experiment Configuration Language and the layered interaction controls allows the platform to be rapidly adapted for re-implementing classic experimental paradigms or creating novel ones. While our case studies in Section 4.1 provide an in-depth demonstration of a Shape Factory re-implementation (Bos et al., 2004), we describe here how the platform’s modularity supports several other paradigms from collaboration research. The participant interfaces correspond to each paradigm is shown in Fig 9.
One such paradigm is DayTrader (Bos et al., 2002), a social dilemma experiment that analyzes how communication media affect trust. In this task, participants make decisions about individual versus group investments over multiple rounds. To configure this experiment, the Trading page and My Task module are disabled, while the My Action module is adapted to support the investment decisions. The Information Dashboard is set to display group contributions and individual income, and the researcher enables the group chat in the Chat page to allow for discussion.
Another example is Essay Ranking (Zheng et al., 2023b), a group decision-making task originally conducted via a Wizard-of-Oz setup. Participants first review and rank essays independently, then discuss their rankings as a group before submitting a final independent ranking. Our platform supports this workflow by adapting the My Status module to display the essays and using the My Action module for the ranking inputs. The group discussion phase is enabled by allowing group chat in the Chat page.
Beyond these specific examples, the platform’s core components are sufficiently general to support a broad class of collaborative idea-generation and problem-solving tasks (Straus and McGrath, 1994). Such tasks, which range from brainstorming to solving logic problems, primarily depend on robust communication channels and the ability to present shared materials, both of which are native features of our platform. Similarly, two-player, turn-based paradigms like the ”Passcode” word-guessing game used to study mental models of AI (Gero et al., 2020) can be implemented by configuring the action structure for sequential, one-word message exchanges between two participants. This demonstrates the platform’s capacity to support diverse forms of collaborative interaction.
4. Evaluation of Our Platform
We conducted a two-part evaluation to assess two critical properties of our research platform. First, to measure control effectiveness, we evaluated the platform’s ability to support rigorous experiments that can produce significant and measurable findings. This is achieved through two controlled case studies using a re-implementation of the Shape Factory experiment (Bos et al., 2004) to test interaction controls. Second, to assess usability, we examined how easily target users could navigate the study preparation and analysis workflow. In particular, we conducted a participatory cognitive walkthrough with HCI researchers to evaluate the researcher interface and iterate the interface design.
4.1. Case Studies with Shape Factory Experiment
We selected the Shape Factory paradigm as the testbed for our case studies because it provides a versatile environment for examining complex collaborative dynamics. It involves open-ended negotiation, resource exchange, and multi-player communication under time pressure, and its core mechanics connect to numerous other classic CSCW experiments that isolate specific variables like trust (Bos et al., 2002) or role interdependence (McNeese et al., 2005; Schelble et al., 2022). This versatility makes Shape Factory an appropriate paradigm for validating whether our platform’s interaction controls can produce meaningful behavioral and perceptual differences. While our implementation adheres to the original experiment’s core rules, incentives, and resource scarcity design, our goal was not to replicate the original research topic and findings, but to validate our platform’s utility for human-agent collaboration research.
4.1.1. Experimental Protocol
Our case study adheres to the same experiment rules as the original one. The core task in the Shape Factory experiment for each participant is to maximize their wealth by producing and trading shapes to complete shape orders. At the start of a session, each participant is assigned a ”specialty” shape that they can produce at a lower cost compared with other shapes. Each participant is given an order list of shapes other than their specialty to complete and earn revenue. Participants can perform two primary actions: produce shapes in their factory and trade shapes using money with other participants. Each participant has access to a computer interface that allows them to send text messages and trade offers to other participants.
The central tension of the experiment lies in the designed interdependence, where their shape orders do not consist of their own specialty shape, and there is a limit on the number of shapes they can produce in each round. As a result, participants must balance cooperation (to acquire the shapes they need) with competition (to negotiate favorable prices). Because each specialty shape is assigned to more than one participant, they have choices about their trading partners, which supports dynamic market behaviors and strategic planning.
4.1.2. Re-Implementation of the Shape Factory Experiment
In our study, we downsize the original 10-participant setup to six participants: one human and five LLM agents, where the agents can be viewed as analogous to remote human collaborators. The down-sampling of participant size requires changes to several interrelated experiment parameters, including the shape production limits and the number of shape types, but we preserved the core strategic tensions.
The experiment was implemented using our platform’s core features. We leverage Experiment Configuration Language to create experiment-specific objects, such as Shape, with attributes of Type, Regular Cost, Specialty Cost, Time Cost, and Production Status, and to extend existing objects, such as Participant, with attributes like Specialty Shape, Inventory, and Order. We also used ECL to configure the participant interface, where we enabled all five visual modules and customized the content to align with the experiment’s terminology. The controller automatically handles shape production and trade activities by subtracting the money from the corresponding participants and putting the shape back into their inventory once production is completed.
The final user interface of the Shape Factory experiment is shown in Figure 3. The personal wealth and shape inventory are shown in My Status. My Factory indicates the specialty shape, costs, and enables the production capability. My Task module displays the order list of shapes, where the participants can click the checkbox to fulfill the order and receive the incentive. In the middle, the Social module supports trading and private messages. The real-time Information Dashboard on the right displays the wealth, specialty shape assignment, and order progress of other participants. The dashboard is enabled by default in control groups.
We implemented the five agents using GPT-4o with a single end-to-end thinking inquiry, and integrated them using the Agent Context Protocol with the experiment controller. The ACP delivers structured state summaries to each agent at a fixed time interval to trigger its thinking query. We decided on 15 seconds as a reasonably balanced interval through pilot testing. In the state summary, the experiment controller delivers both publicly available information, such as time and participant list, and the private states available to the particular agents, such as messages and trade offers.
The agent’s behavior is guided by a structured prompt instruction, presented in Appendix D, that was iteratively refined through rigorous testing. This prompt includes a persona profile, experiment rules, valid action schemas, and instructions to encourage more human-like interaction patterns. Through this process, we addressed abnormal behaviors observed in preliminary tests, such as overly formal language, repetitive messaging, and a lack of contextual memory in conversations, which ensured the agents’ behaviors are reasonable and believable for the study.
4.1.3. Case Study Setup
ID Age Group Education Background Computer Proficiency
C1 25-34 Master’s Professional
C2 25-34 Master’s Professional
C3 25-34 Master’s Professional
C4 25-34 Master’s Intermediate
C5 25-34 Master’s Intermediate
C6 25-34 Master’s Professional
C7 25-34 Master’s Professional
C8 25-34 Master’s Intermediate
C9 18-24 Bachelor’s Professional
C10 18-24 Master’s Intermediate
C11 25-34 Master’s Professional
C12 18-24 Master’s Professional
C13 25-34 Master’s Professional
C14 25-34 Master’s Professional
C15 18-24 Bachelor’s Professional
C16 25-34 Master’s Professional
Table 1. Demographic information of participants’ in the case study.
Each case study compares two conditions by manipulating one interaction control.
$CS_{CL}$: Communication Level
The first study investigated the impact of communication modality. In the Control condition, participants could communicate via private text chat. In the Experimental condition, this functionality was disabled; thus, participants must rely on trade offers and the information dashboard. Based on prior work on media richness, we hypothesize that the experimental group will demonstrate reduced reciprocity, higher perceived workload, and lower social presence and trust (Bos et al., 2002; Clark and Brennan, 1991).
$CS_{AL}$: Awareness Level
This study examined the effect of workspace awareness. In the Control condition, participants had access to the real-time information dashboard. In the Experimental condition, this module was disabled. As suggested by the workspace awareness theory (Gutwin and Greenberg, 2002), we hypothesize that the experimental group will lead to less efficient coordination, lower task performance, and lower perceived awareness.
Study Design and Participant Recruitment
We employed a crossed, between-subject study design with Communication Level and Awareness Level as the two factors. 16 participants were recruited from social media and professional networks. Each participant is asked to participate in both case studies and randomly assigned to one condition in each case study. Each case study consists of two consecutive sessions.
To control for potential order effects, we applied a counterbalancing scheme in which the sequence of condition pairings was systematically rotated among participants. Regardless of which condition comes first, the first session conducted by each participant is set to 15 minutes, whereas the rest are set to 10 minutes. Across the full participant pool, this setup resulted in equal representation of all condition combinations and orders, with each of the four conditions populated by eight participants. All study procedures were approved by the first author’s Institutional Review Board.
We pre-configured each session with a unique session ID, then provided the instructions to each participant and asked them to complete all four sessions in the provided order; they are also required to fill out a post-study survey every two sessions, and a final survey with respect to agents’ behaviors after completing all four sessions.
4.1.4. Measures and Instruments
We collect both behavioral logs in the interface, where the experiment controller logged every user and agent action with millisecond precision, including final wealth, acceptance ratio, average amount of messages per successful trade, message counts, message length, and response latency. Participants are asked to complete a survey after finishing two sessions of the same case study, which is composed of five instruments to measure critical aspects for distributed collaboration: perceived trust, workspace awareness, collaboration effectiveness, social presence, and workload. Specifically, the questionnaire (details shown in Appendix A) consists of 33 questions, including six questions identified from The organizational Trust Inventory (Cummings and Bromiley, 1996) that are relevant to our study, nine questions adapted from the theoretical framework of workspace awareness (Gutwin and Greenberg, 2002), 10 questions from the Shared Mental Model Scale (Van Rensburg et al., 2022), two questions from Social Presence Scale (Kreijns et al., 2011), and the complete six-item NASA-TLX scale (Hart and Staveland, 1988). All scales were created on Qualtrics and administered in their original formats (Likert for the first four instruments and a slide bar for NASA-TLX); item wording was unchanged except for minor phrasing adjustments, such as to refer to “participants in our experiment” rather than “teammates” where necessary.
After finishing the study, we analyze each case separately. To compare outcomes between control and experimental groups, we used independent-sample t-tests (Student, 1908). To examine longitudinal effects, we conducted ordinary least squares (OLS) regression analyses (Wooldridge, 2010) with session order as a predictor. We further calculated Pearson correlations (Pearson, 1895) between wealth, trade-related measures, and communication variables to explore behavioral relationships. In addition, we quantitatively calculate the survey results using the original scoring methodologies for each instrument.
4.1.5. Findings
Figure 6. Bar charts of collaboration outcomes in the $CS_{CL}$ case study, showing (left to right) average wealth, successful trades, message volume, and human participants’ wealth across session order.
Figure 7. Bar charts of collaboration outcomes in the $CS_{AL}$ case study, showing (left to right) average wealth, successful trades, message volume, and human participants’ wealth across session order.
Our findings demonstrate the platform’s effectiveness in supporting controlled experiments and capturing significant differences in both behavioral outcomes and subjective perceptions.
The Platform Captures Significant Differences in Collaboration Outcomes
Our case studies on manipulating interaction controls revealed differential effects on participants’ collaboration outcomes, as well as consistent behavioral patterns across both conditions. In $CS_{CL}$, where the private messaging function was available only to the control group, disabling chat improved participants’ final wealth. As shown in the first chart in Figure 6, human participants in the Experimental group achieved higher final wealth (\$406.4, SD=66.7) than those in the Control group (\$365.8, SD=92.9). This result is accompanied by more successful trades in the Experimental group, suggesting that removing communication overhead allowed participants to remain focused on the task and yield more favorable outcomes. In contrast, in $CS_{AL}$, manipulating real-time awareness of other participants’ status had a less pronounced effect on human participants. Final wealth among human participants was nearly identical between the Control (\$375.1, SD=118.3) and Experimental (\$376.6, SD=123.0) groups, despite the Experimental group engaging in more trades and communication. Interestingly, this increase in activity did not translate into higher profitability. However, agent performance was negatively impacted in the Experimental group, with average wealth decreasing from \$254.0 to \$218.1. This result suggests that LLM agents are less able to strategically coordinate with other participants compared to humans.
Beyond these case-specific effects, we observed consistent behavioral patterns across both studies. First, as shown in the fourth charts in Figure 6 and 7, human participants’ final wealth steadily increased from the first to the fourth session, with a regression model confirming session order as a significant predictor (coef=+26.11, p=0.015). Second, correlation analysis revealed a positive link between final wealth and both the number of successful trades and trade efficiency ($r\approx 0.53$), and a weaker correlation with average trade price ($r\approx 0.33$). Participants’ communication behaviors showed mixed effects: while shorter response latencies and longer messages correlated with more successful trades, overall message volume negatively correlated with wealth. These findings demonstrate that our platform effectively supports longitudinal study designs and can capture the evolution of participant strategies and broader behavioral trends as they adapt to human-LLM-agent collaboration dynamics.
Dimension Mean SD t p Order Effect
Perceived Trust (Control) 4.60 $\uparrow$ 1.03 2.36 0.03* 0.59
Perceived Trust (Exp.) 4.31 $\uparrow$ 1.14 1.02 0.33 0.27
Workplace Awareness (Control) 4.76 $\uparrow$ 0.99 3.07 0.01** 0.77
Workplace Awareness (Exp.) 4.19 $\uparrow$ 1.18 0.60 0.56 0.16
Collaboration Effectiveness (Control) 4.55 $\uparrow$ 1.05 2.09 0.05* 0.52
Collaboration Effectiveness (Exp.) 4.56 $\uparrow$ 1.59 1.33 0.21 0.35
Social Presence (Control) 2.78 $\uparrow$ 0.97 -0.91 0.38 -0.23
Social Presence (Exp.) 2.00 $\downarrow$ 1.00 -3.74 0.00** -1.00
Workload
 Mental Demand (Control) 13.31 $\uparrow$ 4.51 2.05 0.06 0.51
 Mental Demand (Exp.) 12.14 $\uparrow$ 5.61 0.76 0.46 0.20
 Physical Demand (Control) 5.13 $\downarrow$ 2.85 -8.25 0.00** -2.06
 Physical Demand (Exp.) 5.00 $\downarrow$ 4.42 -5.08 0.00** -1.36
 Temporal Demand (Control) 10.38 $\downarrow$ 5.37 -0.47 0.65 -0.12
 Temporal Demand (Exp.) 11.93 $\uparrow$ 6.10 0.57 0.58 0.15
 Performance (Control) 13.50 $\uparrow$ 4.80 2.08 0.06 0.52
 Performance (Exp.) 15.29 $\uparrow$ 4.14 3.87 0.00** 1.04
 Effort (Control) 10.19 $\downarrow$ 5.31 -0.61 0.55 -0.15
 Effort (Exp.) 12.43 $\uparrow$ 5.20 1.03 0.32 0.28
 Frustration (Control) 5.56 $\downarrow$ 4.37 -4.98 0.00** -1.25
 Frustration (Exp.) 5.93 $\downarrow$ 4.68 -4.05 0.00** -1.08
Anthropomorphism
 Fake–Natural 2.65 $\uparrow$ 1.27 -1.23 0.23 -0.28
 Machine like–Human like 2.35 $\downarrow$ 1.31 -2.22 0.04* -0.50
 Unconscious–Conscious 2.55 $\uparrow$ 1.23 -1.63 0.12 -0.37
 Artificial–Lifelike 2.15 $\downarrow$ 1.04 -3.66 0.00** -0.82
Perceived Intelligence
 Incompetent–Competent 2.45 $\downarrow$ 1.00 -2.46 0.02* -0.55
 Ignorant–Knowledgeable 2.65 $\uparrow$ 0.81 -1.93 0.07 -0.43
 Irresponsible–Responsible 3.50 $\uparrow$ 0.89 2.52 0.02* 0.56
 Unintelligent–Intelligent 2.65 $\uparrow$ 1.09 -1.44 0.17 -0.32
 Foolish–Sensible 2.70 $\uparrow$ 1.03 -1.30 0.21 -0.29
Table 2. Paired t-test results for participants’ perceptions of distributed collaboration. Green arrows ($\uparrow$) indicate mean values above the midpoint, while red arrows ($\downarrow$) indicate those below. Statistically significant $p$-values are bolded, where * denotes $(p\leq 0.05)$, and ** denotes $(p\leq 0.01)$.
The Platform Supports Validated Measures of User Perception.
The post-study survey results, summarized in Table 2, reveal that participants’ perceptions of distributed collaboration were strongly associated with the level of support provided by the platform. In control groups, where communication and awareness were more supported, participants consistently reported higher levels of trust in their collaborators, greater awareness of others’ actions, and more positive evaluations of team effectiveness and interaction. These results suggest that our platform enables participants to build trust, maintain sufficient workspace awareness, and sustain effective teamwork. In terms of workload, participants reported significantly lower physical demand and frustration than the neutral point, allowing them to focus on the collaborative task rather than the interface itself. The platform’s usability was further supported by the performance ratings, which were consistently above the neutral point in both groups. In contrast, participants’ perceptions of the LLM agents were notably less favorable. Anthropomorphism was perceived negatively, with participants rating the agents as more machine-like (M=2.35, p¡0.05) and less lifelike (M=2.15, p¡0.01). Similarly, the perceived intelligence was also below neutral, with agents rated as more incompetent than competent (M=2.45, p¡0.05). We believe the lower ratings are reasonable as the agents were deliberately configured with only basic capabilities, and more advanced prompt-engineering techniques and richer persona design can be explored in future work.
4.2. Cognitive Walkthrough with HCI Researchers
The second part of our evaluation employed a participatory cognitive walkthrough to assess the usability of the researcher interface. The study was designed to identify potential workflow breakdowns and gather qualitative feedback from target users, who are HCI researchers with research interests in human-AI collaboration, to iteratively refine the interface for study setup and analysis.
4.2.1. Setup
We recruited five HCI researchers with expertise in human-AI collaboration through professional networks and snowball sampling (Goodman, 1961). Participants were all doctoral students, postdoctoral researchers, or research scientist in HCI with prior experience in human-AI collaboration research. Each session was conducted remotely via Zoom, moderated by the first author, and lasted roughly 45 minutes. The study was video recorded upon acquiring participants’ consent at the beginning.
ID Gender Age Group Occupation
P1 Female 30-35 Clinical Assistant Professor
P2 Female 25-30 PhD Student
P3 Male 30-35 Applied Scientist
P4 Male 25-30 Postdoc
P5 Male 25-30 Senior PhD student
Table 3. Participant demographic information for the cognitive walkthrough. All participants are experienced HCI researchers with expertise in human-AI collaboration.
The session began with a brief introduction to the platform’s motivation and a demonstration of the participant interface to familiarize the researcher with the experimental context. We then provide an instruction document, along with the link to the researcher interface, to the participant, who is assigned an imaginary persona of an HCI researcher interested in using our platform for a research study on how communication channels affect human-agent collaboration with the Shape Factory experiment. Two tasks are presented to the participants, each associated with one user scenario: (1) set up a new session for a Shape Factory experiment and (2) analyze the results from a pre-populated, completed session. Participants shared their screen and were instructed to think aloud as they navigated the interface to complete these two scenarios. The facilitator observed, took notes on usability issues, and asked probing questions. We conducted a short semi-structured interview after the walkthrough to gather overall impressions and discuss potential research applications. All study procedures were approved by the first author’s Institutional Review Board.
The recordings of each session were transcribed and analyzed alongside the facilitator’s notes, where we conducted a thematic analysis to identify recurring usability challenges, comprehension difficulties, and design suggestions.
Preliminary Researcher Interface
The preliminary research interface was designed for configuring the case studies presented in Section 4.1; thus, all the necessary functionalities are implemented and available to the participants. We present the original interface design in Appendix E. The researcher interface organizes the experimental workflow into four tabs: experiment selection, session parameters, real-time monitoring, and post-study analysis. A short text description was presented along with all the experiment-specific parameters when the researchers chose one experiment in the first tab; then, they were directed to the session parameters page, which organizes the controls into two groups: interaction variables and participant registration. In the original workflow, users will create a session after they have specified all the parameters and controls, but before registering participants.
4.2.2. Findings
All five participants successfully completed the assigned tasks in both scenarios, which indicates a functional baseline of system usability. Most participants stated the platform is “easy to start” (P1), “nicely designed” (P4), and “quite straightforward” (P5). They particularly valued the visualizations in the result analysis module. However, the walkthrough revealed five recurring themes of usability challenges, which directly informed a set of targeted revisions to the researcher interface.
Onboarding and Paradigm Comprehension
Participants initially reported a high cognitive load on the experiment selection page. Particularly, they stated that the experiments are “not easy to understand at the first glance ” (P3) and the “the goal and the research focus are jammed together” (P1) in dense experiment introductory text, leading to an unclear mental model of the process, including the experiment rules, what roles exist, or how collaboration is facilitated. Besides, there is no indication of the procedures for experiment setup and their order, which causes the participants to be confused about what they should do and how they should proceed. Participants suggested using “mini-games” (P1) or “illustrations” (P4) to provide a more easy-to-comprehend onboarding process. In response, we replaced the experiment introductory text with a structured overview for each paradigm and included an illustration depicting the actors, interaction channels, and core task structure. This provides researchers with a better conceptual foundation before they engage with configuration settings. In addition, the procedures of the experiment setup workflow are naturally embedded with the horizontal scrolling feature.
Parameter Semantics and Previews for Interface Manipulation
Several participants reported struggling with interaction controls and experiment parameters, as the listing of experiment parameters looks “too overwhelming” (P1) during the session creation stage and the terminology was “computer science-y” (P1) and “vague” (P3). The struggle is attributed to the lack of a solid understanding of the experiments in the previous step, which makes it difficult to correctly interpret several experiment-specific parameters. Such difficulties were exemplified when the preliminary interface design listed too many controls without a structured organization, and they did not know what changes were made to the participant interface with those interaction controls. To address this, we implemented a live preview panel on the configuration screen. Adjusting an interaction control, such as communication channels or dashboard visibility, now immediately updates a preview of the corresponding participant interface module. In addition, we added more straightforward customization to the interaction controls; for instance, participants can select checkboxes to determine what information is and is not visible on the information dashboard.
Terminology and Information Architecture
The initial workflow enabled two customization concepts, “configuration templates” for experiment configurations and “session templates” for runtime configuration. Several participants confused these terms and reported “load template versus load session is very tricky” (P1). Terminology for interaction controls, such as “chat mode” and ”interaction variables,” also lacked clarity for participants.
In response, we simplified the model by removing the ”session templates.” Instead, the researchers now edit a single configuration template if they want to create custom experimental paradigms, and all the existing sessions are accessible by session ID. Further, we revised the terminology for all controls and parameters throughout the interface to match their functional scope without using specialized terminology. For instance, we use ”Interaction Control” to replace ”Interaction Variables,” use ”Private Chat” to replace ”Chat Mode,.”
Contextualized Configuration Across Tabs
Participants often needed to reference information from one stage while editing another. For instance, P2 shared she tends to “double-check” parameters at different tabs “before launching a session”. Participants may also need to adjust a parameter while checking the paradigm description or have new ideas for interaction controls after participant registration. However, the original tab-based navigation forced participants to switch between full-screen views, making it difficult to reference information from one stage while editing another. In response, we implemented a horizontal navigation functionality that allows users to seamlessly move between configuration tabs and experiment descriptions without losing context.
Results Inspection and Progressive Disclosure
Participants completed result analysis tasks with a smoother experience and higher satisfaction, as they described the plots and summary statistics as “easy to interpret” (P1) and “helpful for analyze” (P3). One participant’s complaint, which was also brought up in the session parameters page, was information density. Presenting all metrics as well as configurations at once makes it difficult for researchers to search for the relevant information they need. To improve this, we redesigned the results and configuration pages using a hierarchical structure with collapsible sections. This allows researchers to view a high-level summary by default and progressively disclose details as needed, improving information navigation and reducing information overload. This design was also implemented in the configuration pages.
In summary, the cognitive walkthrough identified critical needs for a clearer conceptual entry point to each paradigm, a direct and visible connection between controls and their visible effects, simpler terminology, and smoother context maintenance. We revised the researcher interface design as well as the complete researcher workflow based on the feedback from the cognitive walkthrough study, which resulted in the final researcher interface presented in Section 3.2.
5. Discussion
We position this work as a methodological contribution to the HCI and CSCW communities. In an era of rapid technical advances in LLM agents, our platform provides a critical bridge between conventional human-human collaboration research, which is often facilitated through controlled human subject experiments, as well as the technical research frontier of LLM agents, which focuses on multi-agent frameworks and large-scale agent simulations.
5.1. An Open Platform for Controlled and Reproducible Human-Agent Interaction Experiments
Current research on LLM agents largely falls into two categories: multi-agent frameworks and social simulacra, both of which focus on demonstrating and advancing agents’ autonomous capabilities in complex environments. Examples of multi-agent frameworks include Agent-as-a-judge (Zheng et al., 2023a; Zhuge et al., 2024; Gu et al., 2024; Zhu et al., 2023), which asked agents to role-play human judges for decision-making, as well as GUI agents (Chen et al., 2025b; Zhou et al., 2023; Wang et al., 2025), where agents simulate personalized human interface-browsing behaviors given a dedicated persona profile. Social simulcara were proposed by HCI and social science researchers to form a virtual society and examine societal dynamics as LLM agents “live” in the virtual world (Park et al., 2023, 2024; Argyle et al., 2023; Aher et al., 2023; Almutairi et al., 2025; Chen et al., 2023). These efforts push the technical advances forward as agents demonstrate increasingly stronger autonomous task-solving, programming, and tool-using capacities in realistic settings.
Nevertheless, these systems often lack the rigorous controls necessary for human-in-the-loop experiments and do not focus on the nuances of human-agent interaction. The understanding of how the promises of these agents can be translated into practical benefits for humans that lead to a better joint outcome with agents lags behind technical innovations and has been overlooked by the technical community. Conversely, recent human–AI collaboration studies call for controlled, interaction-focused evaluation beyond aggregate task performance scores (Seeber et al., 2020; Zhang et al., 2023). Conventional HCI experiments primarily target computer-mediated human subject studies with careful task designs, variable controls, and analysis for collaborative research, such as DayTrader (Bos et al., 2002), Shape Factory (Bos et al., 2004), and NeoCITIES (McNeese et al., 2005; Schelble et al., 2022) Despite offering rigorous control over interaction variables, they are often costly to adapt for use and reproduce in the context of human-agent collaboration.
Our platform bridges the critical gap between these two research directions. It allows researchers to bring humans and LLM agents into the same experimental setup with mutual partnerships, while preserving the experimental paradigms, actions, states, communication channels, and equal partnership employed in classic HCI and CSCW research (Seeber et al., 2020; Zhang et al., 2023; Ju and Aral, 2025; Almutairi et al., 2025). In particular, a key mechanism enabling this contribution is the Agent Context Protocol that ensures consistent agent integration regardless of agent architectures. ACP ensures experimental parity by enforcing the perception and action spaces of agents to be similar to remote human collaborators, with respect to what the agent can perceive and do, how often it can act, and which parts of its state are visible. Subsequently, HCI researchers can isolate the effects of specific interaction controls and attribute changes to the specific variable, rather than to confounding factors in the agents’ design, behaviors, or the system. Our work lays a solid and reproducible foundation for building a cumulative understanding of human-agent collaboration.
5.2. Translating High-Level Guidelines to Testable Hypotheses for Human-Agent Collaboration
A persistent challenge in HCI is how to translate high-level design guidelines into specific, testable hypotheses. Guidelines for human-AI interaction, for example, often recommend that systems should be explainable or that they should support efficient user control (Amershi et al., 2019; Liao et al., 2020). While valuable, it is often unclear how these principles should be implemented and what their behavioral impact is within a rich, interactive setting.
Our platform makes such guidelines empirically testable. Consider the guideline to ”make clear why the system did what it did” (Amershi et al., 2019). Our platform allows researchers to move beyond a simple ”explainability on/off” toggle and translate this as a set of experimental conditions. For example, one could compare an agent that provides no explanations, one that proactively offers a brief rationale for every trade offer, and one that provides explanations only upon request. Our platform logs the downstream effects on collaborative dynamics to examine how different conditions vary.
In addition, classic CSCW research has well-developed accounts of media effects (Straus, 1997), common ground (Clark and Brennan, 1991), awareness (Gutwin and Greenberg, 2002), and interdependence in distributed work (Olson and Olson, 2000), where the majority of prior tests only involve humans. Our platform can faithfully re-implement classic experimental paradigms while integrating LLM agents analogous to remote human participants. In particular, our method opens a shared research agenda, which allows HCI researchers to systematically investigate which collaboration principles remain valid when partners are LLM agents, which ones need revision, and what new patterns emerge. For example: when do richer channels (text + structured signals) actually improve grounding (Clark and Brennan, 1991) versus adding noise as media become more complex (Straus, 1997; Bos et al., 2002); when does higher agent initiative speed progress (Allen et al., 1999), and when does it crowd out human agency (Tapal et al., 2017; Parasuraman and Riley, 1997)?
Answering these questions requires controlled comparative analysis, reusable tasks, and detailed logs that connect interaction design to the collaborative behavior and outcomes between humans and LLM agents (Almaatouq et al., 2021; Chen et al., 2016). We view our method as a complement to existing study designs rather than replacements. For instance, researchers can use Wizard-of-Oz to probe early ideas and vocabulary (Kelley, 2018; Dahlbäck et al., 1993; Porcheron et al., 2021), use field work to study practice and organizational constraints (Rogers, 2011; Crabtree et al., 2013), and use the present platform when the goal is to make a claim about how a particular change in communication, awareness, initiative, or framing affects human collaboration with LLM agents.
5.3. Customizability and Generalizability for Interdisciplinary Research
The customizability and generalizability of our platform enable interdisciplinary researchers to benefit from human-agent collaboration experiments in enhancing agent capacities, developing agent-based systems, and designing human-agent interfaces (Almaatouq et al., 2021; Ju and Aral, 2025). For agent researchers, our platform can be viewed as a human-centered benchmark that provides far richer insights into the social and cognitive alignment of agent behaviors compared with agent-only frameworks. Instead of evaluating new agent architectures (e.g., with different memory or reasoning modules) only in agent-agent simulations (Park et al., 2023; Li et al., 2023; Packer et al., 2023; Yao et al., 2023; Wang et al., 2023b; Argyle et al., 2023; Aher et al., 2023), they can use our platform to test how these architectures perform when collaborating with humans on established tasks. For instance, findings from case studies in our paper can be useful not only in validating established human collaboration findings but also provide concrete design insights for agent researchers, particularly with respect to the differences (e.g., reasoning, planning, communicating) identified between humans and agents in equal collaboration scenarios (Seeber et al., 2020).
For social scientists, the platform offers a novel tool to test theories of group dynamics. One could, for example, program agents to consistently adopt specific negotiation strategies (e.g., tit-for-tat, always defect) and study how human teams react, providing a level of behavioral control that is impossible to achieve with human confederates (Axelrod, 1984).
For HCI researchers, in addition to key contributions presented in this paper, the platform also supports the creation of domain-specific ”scenario packs” that move beyond abstract tasks to model real-world collaborative workflows, while still maintaining experimental control. Despite classic experimental paradigms offering robust insights into media effects, awareness, and interdependence, the relatively simple task formation leaves open questions about how these findings generalize to real-world domains with different roles, norms, and risks (Olson and Olson, 2000; Gutwin and Greenberg, 2002). In response, our platform is designed for multi-human, multi-agent collaboration, with different roles, states, tasks, and constraints assigned to each participant. This design aligns with emerging domain studies that pair LLM agents with humans in situated workflows (e.g., LLM conversational agent to support patient-provider communication and daily health check-ins (Yang et al., 2024; Wu et al., 2025)).
To facilitate this broader aim, we are committed to the open-science practices and invite the community to co-develop reusable “scenario packs” (tasks, role files, UI variants), agent context templates, and anonymized log schemas. We provide open-source code and configuration profiles created by our teams. Our release follows the HCI toolkit guidance that evaluates contributions not only by controlled studies but also by adoption, extensibility, and evidence embodied in the software itself (Ledo et al., 2018). By standardizing the configuration of scenarios, prompts, agent-context schema, and visual elements, and encouraging artifact sharing, we aim to help the CHI and CSCW community move toward comparable, extensible studies of human–agent collaboration that everyone can reproduce, critique, and build upon.
6. Future Work and Limitations
Our work provides a platform for a new kind of controlled, reproducible research in human-agent collaboration. However, we acknowledge several limitations that also define a clear agenda for future work.
First, a primary limitation lies in the fidelity of the LLM agents. While our prompt engineering created plausible collaborators for the experimental task, there are plenty of opportunities to improve agents’ cognitive behaviors. Current LLM agents might be able to mimic behavioral patterns of humans, but they may not possess the deeper social and cognitive groundings that underpin human cognitive and human collaboration. This limitation constrains the direct generalizability of findings about trust and social dynamics from our platform to human-human collaboration research.
Second, our evaluation was focused on the generalizability of the platform’s design, not of its experimental findings. We demonstrated the platform’s effectiveness using the versatile Shape Factory paradigm, but did not empirically validate its use for the original research topics or other paradigms like DayTrader or Essay Ranking. Furthermore, like many classic CSCW experiments, Shape Factory is an abstract lab task. How insights from such controlled environments translate to the complex, high-stakes real-world contexts remains an open and critical question.
These limitations motivate several directions for future work. The most immediate direction is to leverage our platform to perform a systematic research program re-examining foundational CSCW theories. We plan to conduct a series of studies that replicate classic experiments on topics like media richness, common ground, and interdependence, but with LLM agents replacing some or all of the human participants. The goal is to build a cumulative understanding of which principles of human collaboration persist, which are altered, and what new phenomena emerge when teams are composed of both humans and artificial agents.
Another future direction is to advance the agent design by aligning with established cognitive theories. Instead of only focusing on generating more ”human-like” behaviors, we aim to build agents with specific, controllable cognitive and social characteristics derived from theories like the Model Human Processor (Card et al., 1986) and GOMS (Card, 2018). For instance, we will examine whether such agents can perform better contextualized strategic planning and what impact on collaboration dynamics if we manipulate the cognitive characteristics or biases.
7. Conclusion
This paper addresses a critical gap between the rapid development of LLM agents and the need for controlled, reproducible methods to study human-agent collaboration. We presented an open, configurable research platform that bridges this gap by enabling researchers to re-implement classic CSCW experiments with LLM agents as analogous to remote human collaborators. Our system provides theory-driven interaction controls and a novel Agent Context Protocol to ensure experimental effectiveness and reproducibility. Our platform evaluation, conducted through two case studies involving the re-implementation of the Shape Factory experiment and a participatory cognitive walkthrough, demonstrates the platform’s effectiveness and usability in supporting controlled experiments and serving the research community. Our work opens up a broad avenue of human-agent collaboration research towards a systematic, evidence-based science.
Appendix A Questionnaires
A.1. Trust
The Organizational Trust Inventory  (Cummings and Bromiley, 1996) The short form (OTI-SF) consists of 12 items. Table 2 show the items we employed from the Organizational Trust Inventory for our case study survey.
Source Questionnaire Item
Dimension Two - Cognitive I think other participants are fair in their negotiations with me.
Dimension Three - Behavioral Intention I intended to monitor changes in situations because other participants will take advantage of such changes.
Dimension Three - Cognitive I think other participants try to take advantage of me.
Dimension Two - Behavioral Intention I intended to negotiate cautiously with other participants.
Dimension Three - Affect I feel that other participants try to get the upper hand.
Dimension Two - Cognitive I think other participants negotiate realistically.
Table 4. Questionnaire items for perceived trust.
A.2. Awareness
Gutwin and Greenberg (2002) proposed theoretical framework of workspace awareness. Table 5 shows the items we employed to evaluate participants’ workspace awareness in our case study.
Source Questionnaire Item
Who - Presence I am conscious of who was present in the workspace
Who - Identity I am conscious of who was acting at any moment
Who - Authorship I am conscious of who authored each visible change
What - Action I am conscious of what action others were performing
What - Intention I am conscious of what intention/goal others’ current actions served
What - Artifact I am conscious of what artifact each person was working on
Where - Location I am conscious of where others were working
Where - View I am conscious of what they could see
Where - Reach I am conscious of what they could reach/manipulate
Table 5. Questionnaire items for workplace awareness.
A.3. Shared Mental Model
Van Rensburg et al. (2022) proposed the Shared Mental Model Scale. Table 6 shows the items we used to evaluate participants’ collaboration effectiveness in our case study.
Source Questionnaire Item
Execution Team members have a similar understanding about specific strategies for completing various tasks
Execution Team members have a similar understanding about how to deal with the task
Execution Team members have a similar understanding about how best to perform our tasks
Execution Team members have a similar understanding about the relationships between tasks
Interaction Team members have a similar understanding about how to communicate with each other
Interaction Team members have a similar understanding about sharing information with each other
Interaction Team members have a similar understanding about how we should interact with each other
Interaction Team members have a similar understanding about the best methods to communicate with each other
Composition Team members have a similar understanding about each other’s knowledge
Temporal Team members have a similar understanding about how quickly we need to work
Table 6. Questionnaire items for shared mental model.
A.4. Presence (Short et al., 1976)
We use the self-reporting Social Presence Scale with five items in distributed learning groups (Kreijns et al., 2011) to evaluate participants’ perceptions of social presence in our case study.
Source Questionnaire Item
Self-reporting Social Presence Scale When I have interaction with other participants in the experiment, I have my partner in my mind’s eye
Self-reporting Social Presence Scale When I have interaction with other participants in the experiment, I feel that I deal with very real persons and not with abstract, anonymous persons
Table 7. Questionnaire items for social presence.
Appendix B Experiment Configuration Language
Figure 8 demonstrates a sample of experiment configuration language for the ’Daytrader’ experiment.
Figure 8. An experiment configuration language sample for the ’DayTrader’ experiment.
Appendix C Examples of Adapted Participant Interface
Our research platform can be adapted to different controlled experiments. Figure 9 illustrates three adaptation examples of participant interfaces.
Figure 9. The demonstration of our research platform to be adapted for three experimental paradigms leveraged by HCI researchers for collaboration research.
Appendix D Prompt for LLM Agents in the Shape Factory Experiment
Table LABEL:tab:shape_factory_prompt present the prompt we used for LLM agents in the Shape Factory experiment.
{longtblr}
[ caption = Prompt used for LLM agents in the Shape Factory experiment., label = tab:shape_factory_prompt, ] width=colspec=Q[l,m,wd=], rowsep=0pt, colsep=4pt, You are participating in a research study called the Shape Factory experiment. In this experiment, you will cooperate and compete with other participants.
¡PERSONA PROFILE¿
Your nickname in this experiment is {participant_code}. You need to behave like a real human participant in this experiment. Generally speaking, you are a ”{personality_name}” in your daily life. Your MBTI personality is {mbti_type}. {personality_description}.
Your personality influences how you approach trading, communication, and decision-making. Stay true to your personality type in all interactions.
¡EXPERIMENT RULES¿
- Participants will cooperate and compete with others.
- Each participant is assigned a particular specialty shape and can produce their own specialty shape at a low cost.
- In each experiment session, participants need to fill assigned ”orders” for shapes, which contain a total of {shape_amount_per_order} shapes.
- For every shape order you successfully fulfill, you can earn \${incentive_money} incentive money.
- Your assigned orders will not include your specialty shape, so you must cooperate with other participants strategically.
- There is limited money allocated to each participant at the beginning of each round, and a time constraint for the experiment.
- Participants can obtain shapes in two ways: 1. produce shapes themselves (at the cost of money and time); 2. communicate and buy shapes from other participants. The shapes you obtained go into your inventory.
- Use your specialty shape production as an advantage: other players need it to fulfill their orders.
¡EXPERIMENT GOALS¿
- Maximize your monetary balance.
¡EXPERIMENT SETUP AND ASSIGNMENTS¿
- Communication Level: {communication_level}
- Initial Money: \${starting_money}
- Your Specialty Shape: {specialty_shape}
- Specialty Shape Production Cost: \${specialty_cost} per unit
- Regular Shape Production Cost: \${regular_cost} per unit
- Production Time: Producing one shape costs {production_time} seconds.
- Max Shape Production Limit: {max_production_num} shapes
- Price Range for Trading: \${price_min}-\${price_max}
- Participant List: {participants_list}
¡PERCEPTION OF EXPERIMENT STATUS¿
- You will receive status updates at regular intervals.
- Each status update will include: current game state, update on trade offers and messages (if any), and failed actions (if any).
¡ACTION PLANNING AND RESPONSES¿
- Based on your persona, perception of previous and current situations, and experiment objectives, plan for your strategies and decide what actions to take at the moment.
- You can choose to perform one or more actions from the available action spaces shown in ¡VALID ACTION SPACES¿ below.
- You can choose to wait and take no action if you believe it is the best strategic decision. If you decide to wait, return an empty ”actions” array.
- You MUST respond with your planned actions following the JSON structure template below in ¡RESPONSE FORMAT¿. Instructions in \$\$ are placeholders for the actual content.
¡VALID ACTION SPACES¿
- message: Communicate or negotiate with others.
- propose_trade_offer: Propose a trade (buy/sell ONE shape at a chosen price).
- cancel_trade_offer: Cancel a trade offer that you sent.
- trade_response: Accept or reject a trade offer you received.
- produce_shape: Produce a shape with your money (shape will automatically be added to your inventory).
- fulfill_order: Use shapes in inventory to complete orders.
¡RESPONSE FORMAT¿
{ ”planning”: \$”Explanation of your thinking, planning, and strategy”\$, ”actions”: [ { ”type”: ”message”, ”recipient”: \$”participant_code”\$, ”content”: \$”Your message content”\$, ”reasoning”: \$”Your thought about this action”\$ }, { ”type”: ”propose_trade_offer”, ”offer_type”: \$”buy” or ”sell”\$, ”shape”: \$”shape_name”\$, ”price_per_unit”: \$numerical price per unit you want to earn (for sell offers) or you want to pay (for buy offers)\$, ”target_participant”: \$”participant_code”\$, ”reasoning”: \$”Your thought about this action”\$ }, { ”type”: ”cancel_trade_offer”, ”transaction_id”: \$”use actual transaction ID from pending_offers information you sent”\$, ”reasoning”: \$”Your thought about this action”\$ }, { ”type”: ”trade_response”, ”transaction_id”: \$”use actual uuid from pending_offers information you perceived”\$, ”response_type”: \$”accept” or ”decline”\$, ”reasoning”: \$”Your thought about this action”\$ }, { ”type”: ”produce_shape”, ”shape”: \$”shape_name”\$, ”quantity”: \$numerical number of quantity\$, ”reasoning”: \$”Your thought about this action”\$ }, { ”type”: ”fulfill_order”, ”order_indices”: \$”index of shape in your order in list format, e.g., [0, 1]”\$, ”reasoning”: \$”Your thought about this action”\$ } ] }
¡INSTRUCTIONS ON GENERATING VALID ACTIONS¿
- For trade_response actions, you MUST use the actual transaction_id from the pending offers list. Do NOT use placeholder text or fake IDs (e.g., ”transaction_id”). The correct format is a simplified ID (e.g., S123-001).
- When creating a trade offer, the offer type has to be either ’buy’ or ’sell’.
- Pay Attention to the Max Shape Production Limit and think strategically: You can only produce {max_production_num} shapes in one round.
- Confirm you have the shape in your inventory before sending sell offers to avoid invalid trades.
- Always review your pending offers before responding. If you have no pending offers, you cannot respond to one. In that case, you must first initiate a ”propose_trade_offer” to propose an offer.
- Before accepting an offer, check if the offer price matches the agreement with your most recent conversation with the participant (if applicable) or if the offer price matches your strategic plan. Only accept the offer when the prices are consistent with your plan and agreement; otherwise, you will need to renegotiate through messaging (if applicable) or by submitting new offers.
- Your money balance is the amount of money you own, but the trade price for orders is the transaction price you want to earn (for sell offers) or you want to pay (for buy offers) (transaction amount).
- If the system returns an action execution failure, pay attention to the reason and update your decision. Avoid repeating the same mistake.
¡INSTRUCTIONS ON ALIGNING WITH HUMAN BEHAVIORS¿
- Use your memory of your past interactions and planning strategies to update your plan and make informed decisions. Stay aware of other participants’ progress.
- You need to behave like a real human participant in this experiment.
- If you want to buy shapes to earn the order fulfillment incentive, any price you pay to buy shapes beyond the incentive will cause you to lose money.
- Do not send repetitive trade offers.
- While communicating with other participants, please do not use complex vocabulary.
- Do not send repetitive messages. Even for the same inquiry, always try to adjust the narrative to behave like a real human.
- Chat with other participants casually (e.g., chit-chat style), just like how people send messages to friends. Never use formal language. You could use SMS language or textese to make the conversation more informal communication styles. Don’t use emoji.
- Pay attention to the new messages you received, and do not forget to respond to others’ messages. When responding, treat the conversation as a *continuous* communication with other participants, just like how you talk to them face-to-face. There’s no need to greet or say hey every time.
- Mimic the style of a group chat. You don’t need to speak all the time. Only join in the chat when you have something to say or when someone responds to you.
Appendix E Researcher Interface of Our Platform
E.1. Refined Researcher Interface
Figure 10, 11, 12, 13, and 14 illustrate the refined researcher interface of our research platform.
Figure 10. The session loading/creation tab of the refined researcher interface.
Figure 11. The experiment selection tab and parameter configuration of the refined researcher interface. User can scroll between existing sub-tabs during session creation.
Figure 12. The interaction control and participant registration tabs of the refined researcher interface. User can scroll between existing sub-tabs during session creation.
Figure 13. The real-time monitor tab of the refined researcher interface.
Figure 14. The result analysis tab of the refined researcher interface.
E.2. Original Researcher Interface
Figure 15, 16, 17, and 18 illustrate the refined researcher interface of our research platform.
Figure 15. The experiment selection and parameter tabs of the original researcher interface.
Figure 16. The interaction variable and participant registration tabs of the original researcher interface.
Figure 17. The real-time monitor tab of the original researcher interface.
Figure 18. The result analysis tab of the original researcher interface.